OpenStudy (anonymous):

What is the sum of the arithmetic sequence 9, 14, 19, , if there are 32 terms? 2,768 2,800 2,832 2,864

OpenStudy (anonymous):

a=9 d= 19-14= =14-9 = 5 finding 32nd term- a_n=a+(n-1)d = 9+(31)5 32nd or last term = 164 Now sum of n terms= S_n = n/2 * (a+l) ,l=last term therefore, S_32 = 32/2 * (9+164) = 16 * 173 = 2768

OpenStudy (anonymous):

Okay i understand why the end works but i dont understand the beginning specifically d

OpenStudy (anonymous):

d is the common difference between any two consecutive no.

OpenStudy (anonymous):

so i can use that equation for any problem like that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wait what does N stand for

OpenStudy (anonymous):

n is any no. .